The product of two numbers is 8÷3.If one number is 3÷5,find their sum.
Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)



I already answered this one.

John

My calculator said it, I believe it, that settles it


Answer by Alan3354(69443)

The product of two numbers is 8÷3.If one number is 3÷5,find their sum.
========
m*n = 8/3
m = 3/5
-----
(3/5)n = 8/3
3n/5 = 8/3
9n/5 = 8
9n = 40
n = 40/9
-------
3/5 + 40/9 = (3*9 + 5*40)/45
= (27 + 200)/45
sum = 227/45